# Finely Stratified Rerandomization Designs

### v2

@@ 5.1 Minimax Rerandomization
Since $\gamma_{0}$ is unknown at design-time, we instead take a minimax approach that incorporates prior information about the coefficient $\gamma_{0}$. For information set $B\subseteq\mathbb{R}^{d_{h}}$ specified by the researcher, consider rerandomizing until the worst case in-sample correlation between treatments and covariates is small enough,
$\sup_{\gamma\in B}|\gamma^{\prime}\mathcal{I}_{n}|=\sup_{\gamma\in B}\left|\frac{\operatorname{Cov}_{n}(\gamma^{\prime}h_{i},D_{i})}{\operatorname{Var}_{n}(D_{i})}\right|\leq\epsilon.$ (5.1)
Equivalently, we rerandomize until $p_{B}(\mathcal{I}_{n})\leq\epsilon$ for the convex penalty function $p_{B}(x)=\sup_{\gamma\in B}|\gamma^{\prime}x|$. This significantly generalizes the commonly used quadratic penalty function $p(x)=x^{\prime}\Sigma x$ (e.g. Schindl and Branson, (2024)). Our next result shows that Equation 5.1 is a linear rerandomization, characterizing the implicit acceptance region $A$.
@@
The criterion $p_{B}(\mathcal{I}_{n})\leq\epsilon$ $\iff$ $\mathcal{I}_{n}\in A_{0}$ for $A_{0}=\epsilon B^{\circ}$ with $B^{\circ}=\{x:\sup_{\gamma\in B}|\gamma^{\prime}x|\leq 1\}\subseteq\mathbb{R}^{d_{h}}$, the absolute polar of $B$. The set $A_{0}$ is symmetric and convex. If $B$ is bounded, $A_{0}$ is closed and has non-empty interior. [^23] Also if $\operatorname{int}B\not=\emptyset$ then $A_{0}$ is bounded. See Aliprantis and Border, (2006) for more on polar sets.
@@
Note that since $A_{0}$ is symmetric, the discussion after Theorem 3.5 implies that the asymptotic distribution of $\widehat{\theta}$ under the design in Equation 5.1 is centered at zero. We let $B$ be totally bounded in what follows. The proposition shows that in this case $A_{0}$ has non-empty interior, satisfying the conditions of Assumption 3.1.
@@
We showed above that $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})|W_{1:n}\Rightarrow L_{0}$ for $L_{0}=\mathcal{N}(0,V(\gamma_{0}))+\gamma_{0}^{\prime}Z_{hA}$. Since $\gamma_{0}$ is unknown at design time, define a family $\{L_{\gamma,A}:\gamma\in\mathbb{R}^{d_{h}},A\subseteq\mathbb{R}^{d_{h}}\}$ of possible limiting distributions of $\widehat{\theta}$, with each $L_{\gamma,A}=\mathcal{N}(0,V(\gamma))+\gamma^{\prime}Z_{hA}$ a sum of independent RV’s as above. For any distribution in this family, the conditional asymptotic bias of $\widehat{\theta}$ given realized covariate imbalances $Z_{hA}$ is $\operatorname{bias}(L_{\gamma,A}|Z_{hA})\equiv E[L_{\gamma,A}|Z_{hA}]$. Our main result shows that the polar acceptance region $A_{0}=\epsilon B^{\circ}$ maximizes rerandomization acceptance probability $P(Z_{h}\in A)$, subject to a strict constraint on asymptotic conditional bias, uniformly over all limiting distributions consistent with the prior information $\gamma_{0}\in B$.
@@
$A_{0}=\operatorname*{argmax}_{A\subseteq\mathbb{R}^{d_{h}}}P(Z_{h}\in A)\quad\text{s.t.}\quad\,\sup_{\gamma\in B}|\operatorname{bias}(L_{\gamma,A}|Z_{hA})|\leq\epsilon.$ (5.2)
@@
The expected number of independent draws of $Z_{h}$ until $Z_{h}\in A$ is $P(Z_{h}\in A)^{-1}$. Then, equivalently, Equation 5.2 shows that the acceptance region $A_{0}$ minimizes the (asymptotic) computational cost of rerandomization, subject to a strict guarantee on statistical performance. From the final statement of the theorem, if the prior information set $B$ is well-specified ($\gamma_{0}\in B$), setting $A_{0}=\epsilon B^{\circ}$ bounds the magnitude of the conditional asymptotic bias $E[L_{0}|Z_{hA_{0}}]$ of the GMM estimator $\widehat{\theta}$ below $\epsilon$. In particular, this implies that the variance $\operatorname{Var}(L_{0})$ of the asymptotic distribution $\sqrt{n}(\widehat{\theta}-\theta_{n})\Rightarrow L_{0}=\mathcal{N}(0,V_{a})+\gamma_{0}^{\prime}Z_{hA_{0}}$ is within $\epsilon^{2}$ of the optimal partially linear variance $V_{a}$ in Equation 3.7.
@@
Before continuing, we briefly note an interesting interpretation of the design in Equation 5.1. For distributions $P,Q$ and a function class $\mathcal{F}$, the integral probability metric is $\rho(P,Q;\mathcal{F})\equiv\sup_{f\in\mathcal{F}}|E_{P}[f(X)]-E_{Q}[f(X)]|$. [^25] The pseudometric $\rho$ is also referred to as the maximum mean discrepancy. This is a commonly used statistic in two-sample testing, see e.g. Gretton et al., (2008). Let $\mathcal{F}_{B}=\{\gamma^{\prime}h:\gamma\in B\}$ and define empirical distributions $\widehat{P}_{d}=(h_{i}|D_{i}=d)$ for $d=0,1$. We have
$\sup_{\gamma\in B}|\gamma^{\prime}\mathcal{I}_{n}|\leq\epsilon\iff\sqrt{n}\rho(\widehat{P}_{1},\widehat{P}_{0};\mathcal{F}_{B})\leq\epsilon.$
@@
The minimax design rerandomizes until covariates $h$ are balanced according to $\rho(\widehat{P}_{1},\widehat{P}_{0};\mathcal{F}_{B})$, a distance between covariate distributions that is only sensitive to the projections $\gamma^{\prime}h$ that actually matter for estimating $\theta_{n}=\operatorname{SATE}$. By doing so, we maximize acceptance probability subject to the statistical guarantee in Theorem 5.2.
5.2 Specifying Prior Information
Without pilot data, we must use introspection to choose the prior information set $B$ containing $\gamma_{0}$. Intuitively, $\gamma_{0}$ parameterizes how much the average outcome level $\bar{Y}$ can change for a unit change in $h$, holding $\psi$ fixed. If $t_{0}(\psi)=t^{\prime}\psi$ happens to be linear, then $\bar{Y}=c+\gamma_{0}^{\prime}h+t^{\prime}\psi+e$ and $\gamma_{0}$ is an OLS coefficient. The following examples provide some reasonable prior information specifications and their associated acceptance regions, using a general characterization of acceptance regions $A_{0}=\epsilon B^{\circ}$ in Lemma 5.6 below.
Example 5.4 (Rectangle).
One specification is to assume $\gamma_{0j}\in[l_{j},u_{j}]$ for each $1\leq j\leq d_{h}$, setting $B=\prod_{j=1}^{d_{h}}[l_{j},u_{j}]$. This allows for sign constraints, e.g. $0\leq\gamma_{0j}\leq m$ for some $j$ and $-m\leq\gamma_{0j}\leq 0$ for others. Lemma 5.6 below shows that if $B=\prod_{j=1}^{d_{h}}[l_{j},u_{j}]$, then $A_{0}=\epsilon B^{\circ}=\{x:|x^{\prime}l+x^{\prime}u|+\sum_{j}|x_{j}|u_{j}-|x_{j}|l_{j}\leq 2\epsilon\}$, where $l=(l_{j})_{j}$ and $u=(u_{j})_{j}$. An example is shown in Figure 2. Note that the region $A_{0}$ is conservative in the direction of the set $B=[1,2]\times[1,3/2]$, rejecting covariate imbalances that are too closely aligned with adverse coefficient values $\gamma_{0}\in B$. The region $A_{0}$ is more lenient in directions approximately orthogonal to $B$.
Example 5.5 (Ellipse).
Another natural specification is to set $B=\bar{\gamma}+B_{2}(0,m)$, for an uncertainty parameter $m$ and guess $\gamma_{0}\approx\bar{\gamma}$. By Lemma 5.6, $A_{0}=\{x:|x^{\prime}\bar{\gamma}|+m|x|_{2}\leq\epsilon\}$. More generally, if $B=\bar{\gamma}+\Sigma B_{2}(0,1)$ for a positive-definite matrix $\Sigma$, then $A_{0}=\{x:|x^{\prime}\bar{\gamma}|+|\Sigma x|_{2}\leq\epsilon\}$. One application of this specification is when $B$ is a Wald confidence region constructed using pilot data, as discussed below.
Refer to caption Refer to caption
Figure 2: Prior information $B$ and $A_{0}=\epsilon B^{\circ}$ for Examples 5.4 and 5.5.
The following lemma provides a useful characterization of the acceptance region $A_{0}=\epsilon B^{\circ}$ from Theorem 5.2 for a large family of prior information set specifications. To state the lemma, recall that $|x|_{p}=(\sum_{j}|x_{j}|^{p})^{1/p}$ for $p\in[1,\infty)$ and $|x|_{\infty}=\max_{j}|x_{j}|$. For $p\in[1,\infty]$, denote $B_{p}(0,1)=\{x:|x|_{p}\leq 1\}$.
Lemma 5.6 (Acceptance Regions).
For $p\in[1,\infty]$, let $11/p+1/q=$ 1, setting $q=1$ if $p=\infty$ and vice-versa. Suppose $B=b+\Sigma B_{p}(0,1)$, for $b\in\mathbb{R}^{d_{h}}$ and $\Sigma$ invertible. Then $A_{0}=\{x:|x^{\prime}b|+|\Sigma^{\prime}x|_{q}\leq\epsilon\}$.
5.3 Using Pilot Data
@@
Next, we discuss an alternative strategy that uses pilot data to specify the set $B$. Suppose we have access to $\mathcal{D}_{pilot}\perp\!\!\!\perp(W_{1:n},D_{1:n})$ of size $m$. Suppose $\sqrt{m}(\widehat{\gamma}_{pilot}-\gamma_{0})\approx\mathcal{N}(0,\widehat{\Sigma}_{pilot})$ for some pilot estimator $\widehat{\gamma}_{pilot}$, discussed below. Consider forming the Wald region $\widehat{B}_{pilot}=\{\gamma:m(\widehat{\gamma}_{pilot}-\gamma)^{\prime}\widehat{\Sigma}_{pilot}^{-1}(\widehat{\gamma}_{pilot}-\gamma)\leq c_{\alpha}\}$ using critical value $P(\chi^{2}_{d_{h}}\leq c_{\alpha})=1-\alpha$ for $\alpha\in(0,1)$. Equivalently, one can write this Wald region as
@@
By Example 5.5, using $\widehat{B}_{pilot}$ as a prior information set gives acceptance region
@@
Note that the acceptance region $\widehat{A}_{pilot}$ expands as the pilot size $m$ is larger. This reflects smaller uncertainty about the true parameter $\gamma_{0}$, and thus less adversarial worst case imbalance $\sup_{\gamma\in\widehat{B}_{pilot}}|\gamma^{\prime}\mathcal{I}_{n}|$. Conversely, $\widehat{A}_{pilot}$ shrinks as the confidence parameter $\alpha$ and the scale of the variance estimate $\widehat{\Sigma}_{pilot}$ increases, reflecting greater uncertainty and a more conservative approach to covariate balances. Our next result shows that rerandomization with acceptance region $\widehat{A}_{pilot}$ controls the variance of the residual imbalance $R_{A}=\gamma_{0}^{\prime}Z|Z\in\widehat{A}_{pilot}$ with high probability, marginally over the realizations of the pilot data. The result is an immediate consequence of Theorem 3.5 and Theorem 5.2.
@@
Formally, the pilot estimate of $\gamma_{0}$ and Wald region could be constructed as in Robinson, (1988). In practice, a simple approach suggested by the theory is to let $\widehat{\gamma}_{pilot},\widehat{\Sigma}_{pilot}$ be point and variance estimators from the regression $Y_{T}\sim 1+h+\psi$, for the “tyranny of the minority” (Lin, (2013)) outcomes $Y_{T}=(1-p)DY/p+p(1-D)Y/(1-p)$, noting that $E[Y_{T}|W]=(1-p)Y(1)+pY(0)=\bar{Y}$.
General Parameters. For completeness, we extend the preceding work to general parameters $\theta_{n}$ as in Definition 2.5. Let $\Pi a(W,\theta_{0})$ be the assignment influence function. As in Equation 3.7, consider the partially linear decomposition
$\Pi a(W,\theta_{0})=\gamma_{0}^{\prime}h+t_{0}(\psi)+e,\quad\;E[e|\psi]=0,\;E[eh]=0.$
Note that $e\in\mathbb{R}^{d_{\theta}}$ and $E[e|\psi]=0$ is interpreted componentwise. Consider prior information sets $B_{j}$ for each $\gamma_{0}^{j}$ with $1\leq j\leq d_{\theta}$, where $\gamma_{0}^{j}\in\mathbb{R}^{d_{h}}$ is the $j$th column of $\gamma_{0}$. The final result of this section bounds the asymptotic imbalance term $R_{A}$ if all these prior information sets are well specified.
Theorem 5.8.
Let $D_{1:n}$ as in Definition 2.1 with $A=\cap_{j=1}^{d_{\theta}}\epsilon B^{\circ}_{j}$. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R_{A}$, as defined in Theorem 3.5. If $\gamma_{0}^{j}\in B_{j}$ $\forall j$, then $\max_{j=1}^{d_{\theta}}\operatorname{Var}((R_{A})_{jj})\leq\epsilon^{2}$.
Note that by construction the conservative acceptance region $A=\cap_{j=1}^{d_{\theta}}\epsilon B^{\circ}_{j}$ is symmetric and convex.

### v3

@@ 5.1 Minimax Rerandomization
Since $\gamma_{0}$ is unknown at design-time, we instead take a minimax approach that incorporates prior information about the coefficient $\gamma_{0}$. For belief set $B\subseteq\mathbb{R}^{d_{h}}$ specified by the researcher, consider rerandomizing until the worst case imbalance consistent with $B$ is small enough,
$\sup_{\gamma\in B}|\gamma^{\prime}\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})|\leq\epsilon.$ (5.2)
Equivalently, for imbalance $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$ we rerandomize until $p_{B}(\mathcal{I}_{n})\leq\epsilon$ for the convex penalty function $p_{B}(x)=\sup_{\gamma\in B}|\gamma^{\prime}x|$. This significantly generalizes the quadratic imbalance penalty $p(x)=x^{\prime}\operatorname{Var}(h)^{-1}x$ implicitly used by Mahalanobis rerandomization (Example 2.4). Our next result shows that Equation 5.2 is a linear rerandomization design, characterizing the induced acceptance region $A$.
Refer to caption Refer to caption
Figure 2: Prior information $B$ and $A_{0}=\epsilon B^{\circ}$ for Example 5.2.
@@
The criterion $p_{B}(\mathcal{I}_{n})\leq\epsilon$ $\iff$ $\mathcal{I}_{n}\in A_{0}$ for $A_{0}=\epsilon B^{\circ}$ with $B^{\circ}=\{x:\sup_{\gamma\in B}|\gamma^{\prime}x|\leq 1\}\subseteq\mathbb{R}^{d_{h}}$, the absolute polar set of $B$. The set $A_{0}$ is symmetric and convex. If $B$ is bounded, $A_{0}$ is closed and has non-empty interior. [^27] Also if $\operatorname{int}B\not=\emptyset$ then $A_{0}$ is bounded. See Aliprantis and Border, (2006) for more on polar sets.
@@
Note that since $A_{0}$ is symmetric, the discussion after Theorem 3.5 implies that the asymptotic distribution of $\widehat{\theta}$ under the design in Equation 5.2 is centered at zero. We let $B$ be totally bounded in what follows. The proposition shows that in this case $A_{0}$ is a “nice” set: symmetric, convex, and with non-empty interior, satisfying the conditions of Assumption 3.1.
Dimension Reduction. The oracle region $A=\{x:|\gamma_{0}^{\prime}x|\leq\epsilon\}$ reduced the rerandomization problem to one dimension for arbitrary $\dim(h)$. Similarly, the minimax acceptance region $A_{0}=\epsilon B^{\circ}$ can be viewed as a “soft” form of dimension reduction. To see this, note that the region $A_{0}$ is very stringent about imbalances $\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$ aligned with our belief set $B$, but can allow large imbalances in directions approximately orthogonal to $B$, effectively downweighting these directions in the space of covariates $h$. This effect can be seen in the following example, depicted in Figure 2.
Example 5.2 (Ball).
One natural belief specification is to set $B=\bar{\gamma}+B_{2}(0,u)$, for an uncertainty parameter $u$ and a priori coefficient guess $\bar{\gamma}\approx\gamma_{0}$. Lemma 5.3 below derives the corresponding acceptance region $A_{0}=\{x:|x^{\prime}\bar{\gamma}|+u|x|_{2}\leq\epsilon\}$. For small $u$, acceptance region $A_{0}$ mimics the oracle, allowing very large imbalances $\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$ as long as $\bar{\gamma}^{\prime}\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})\approx 0$. For larger $u$, $A_{0}$ penalizes imbalances in all directions, with a slight extra penalty for being aligned with the coefficient guess $\bar{\gamma}$. This provides a sliding scale of dimension reduction, allowing us to continuously transition between full-dimensional $h$ and one-dimensional $\bar{\gamma}^{\prime}h$ depending on the uncertainty level $u$.
More generally, the following lemma provides a useful characterization of the acceptance region $A_{0}=\epsilon B^{\circ}$ from Theorem 5.5 for a large family of specifications of the belief set $B$. To state the lemma, recall that $|x|_{p}=(\sum_{j}|x_{j}|^{p})^{1/p}$ for $p\in[1,\infty)$ and $|x|_{\infty}=\max_{j}|x_{j}|$. For $p\in[1,\infty]$, denote $B_{p}(0,1)=\{x:|x|_{p}\leq 1\}$.
Lemma 5.3 (Belief Specification).
For $p\in[1,\infty]$, let $11/p+1/q=$ 1. Suppose beliefs $B=\bar{\gamma}+UB_{p}(0,1)$, for $\bar{\gamma}\in\mathbb{R}^{d_{h}}$ and $U$ invertible. Then $A_{0}=\{x:|x^{\prime}\bar{\gamma}|+|U^{\prime}x|_{q}\leq\epsilon\}$.
Example 5.4 (Rectangle).
Assume $\gamma_{0j}\in[a_{j},b_{j}]$ for each $1\leq j\leq d_{h}$, so $B=\prod_{j=1}^{d_{h}}[a_{j},b_{j}]$. This allows for sign and magnitude constraints, e.g. $0\leq\gamma_{0j}\leq m$ for some $j$ and $-m\leq\gamma_{0j}\leq 0$ for others. Lemma 5.3 shows that the acceptance region has form $A_{0}=\epsilon B^{\circ}=\{x:|x^{\prime}(a+b)/2|+(1/2)\sum_{j}|x_{j}|b_{j}-|x_{j}|a_{j}\leq\epsilon\}$, for $a=(a_{j})_{j}$, $b=(b_{j})_{j}$.
5.2 Minimizing Computational Cost
Intuitively, by ignoring imbalances $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$ approximately orthogonal to our beliefs $B$, we can “stretch” the acceptance region $A_{0}$ in directions unlikely to cause large estimation errors, increasing the probability of acceptance $P(\mathcal{I}_{n}\in A)$. Since the expected number of independent randomizations until acceptance is $P(\mathcal{I}_{n}\in A)^{-1}$, we can view this as minimizing the computational cost of rerandomization, subject to a bound on estimation error. This intuition is formalized in Theorem 5.5 below. To state the theorem, we first define the family of possible limiting distributions of $\widehat{\theta}$ consistent with our beliefs $\gamma_{0}\in B$ and choice of acceptance region $A\subseteq\mathbb{R}^{d_{h}}$.
@@ 5.2 Minimizing Computational Cost
Limiting Distributions. We showed above that $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})|W_{1:n}\Rightarrow L_{0}$ for $L_{0}=\mathcal{N}(0,V(\gamma_{0}))+\gamma_{0}^{\prime}Z_{hA}$. Since $\gamma_{0}$ is unknown, define a family of possible limiting distributions of $\widehat{\theta}$ by $\mathcal{L}_{B}=\{L_{\gamma A}:\gamma\in B,A\subseteq\mathbb{R}^{d_{h}}\}$, with each $L_{\gamma A}=\mathcal{N}(0,V(\gamma))+\gamma^{\prime}Z_{hA}$ a sum of independent RV’s. For any distribution in this family, the conditional asymptotic bias of $\widehat{\theta}$ given realized covariate imbalances $Z_{hA}$ is $\operatorname{bias}(L_{\gamma A}|Z_{hA})\equiv E[L_{\gamma A}|Z_{hA}]$. Our main result shows that the polar acceptance region $A_{0}=\epsilon B^{\circ}$ minimizes asymptotic computational cost $P(Z_{h}\in A)^{-1}$, subject to a strict constraint on conditional bias, uniformly over all limiting distributions consistent with our beliefs.
@@
$A_{0}=\operatorname*{argmin}_{A\subseteq\mathbb{R}^{d_{h}}}P(Z_{h}\in A)^{-1}\quad\text{s.t.}\quad\,\sup_{\gamma\in B}|\operatorname{bias}(L_{\gamma A}|Z_{hA})|\leq\epsilon.$ (5.3)
@@
The final statement of the theorem shows that if $B$ is well-specified ($\gamma_{0}\in B$), setting $A_{0}=\epsilon B^{\circ}$ bounds the magnitude of the conditional asymptotic bias $E[L_{0}|Z_{hA_{0}}]$ of the GMM estimator $\widehat{\theta}$ above by $\epsilon$. By the law of total variance, this implies that the variance $\operatorname{Var}(L_{0})$ of the asymptotic distribution $\sqrt{n}(\widehat{\theta}-\theta_{n})\Rightarrow L_{0}=\mathcal{N}(0,V_{a})+\gamma_{0}^{\prime}Z_{hA_{0}}$ is within $\epsilon^{2}$ of the optimal partially linear variance $V_{a}$ in Equation 3.6.
Results closely related to Theorem 5.5 can also be found in the previous work of Liu et al., (2023), who derive optimal Mahalanobis-style completely rerandomized designs under a Bayesian criterion, with Gaussian prior on $\gamma_{0}$.
@@
We briefly note another interesting interpretation of the design in Equation 5.2. For distributions $P,Q$ and a function class $\mathcal{F}$, the integral probability metric is a pseudo-distance between distributions defined by $\rho(P,Q;\mathcal{F})\equiv\sup_{f\in\mathcal{F}}|E_{P}[f(X)]-E_{Q}[f(X)]|$. [^29] The pseudometric $\rho$ is also referred to as the maximum mean discrepancy. This is a commonly used statistic in two-sample testing, see e.g. Gretton et al., (2008). Let function class $\mathcal{F}_{B}=\{\gamma^{\prime}h:\gamma\in B\}$ and let $\widehat{P}_{d}$ denote the empirical distribution of $h_{i}|D_{i}=d$ for $d=0,1$. Then we have
$\sup_{\gamma\in B}|\gamma^{\prime}\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})|\leq\epsilon\iff\sqrt{n}\rho(\widehat{P}_{1},\widehat{P}_{0};\mathcal{F}_{B})\leq\epsilon.$
@@
This shows that the minimax design rerandomizes until within-arm empirical distribution of covariates $h$ are balanced according to $\rho(\widehat{P}_{1},\widehat{P}_{0};\mathcal{F}_{B})$, a distance between $h_{i}|D_{i}=1$ and $h_{i}|D_{i}=0$ only sensitive to the projections $\gamma^{\prime}h$ with $\gamma\in B$ that we believe matter for estimating $\theta_{n}=\operatorname{SATE}$. By doing so, we maximize the size of the acceptance region (and acceptance probability) subject to the statistical guarantee in Theorem 5.5.
5.3 Beliefs From Pilot Data
@@
Next, we discuss an alternative strategy that uses pilot data to specify the set $B$ in a data-driven way. Suppose we have access to $\mathcal{D}_{pilot}\perp\!\!\!\perp(W_{1:n},D_{1:n})$ of size $m$. Suppose $\sqrt{m}(\widehat{\gamma}_{pilot}-\gamma_{0})\approx\mathcal{N}(0,\widehat{\Sigma}_{pilot})$ for some pilot estimator $\widehat{\gamma}_{pilot}$, discussed below. Consider forming the Wald region $\widehat{B}_{pilot}=\{\gamma:m(\widehat{\gamma}_{pilot}-\gamma)^{\prime}\widehat{\Sigma}_{pilot}^{-1}(\widehat{\gamma}_{pilot}-\gamma)\leq c_{\alpha}\}$ using critical value $P(\chi^{2}_{d_{h}}\leq c_{\alpha})=1-\alpha$ for $\alpha\in(0,1)$. Equivalently, one can write this Wald region as
@@
Viewing this $1-\alpha$ confidence region as a belief set, Lemma 5.3 above implies that the corresponding minimax acceptance region is
@@
Note that the acceptance region $\widehat{A}_{pilot}$ expands as the pilot size $m$ is larger. This reflects smaller uncertainty about the true parameter $\gamma_{0}$, and thus less adversarial worst case imbalance $\sup_{\gamma\in\widehat{B}_{pilot}}|\gamma^{\prime}\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})|$. Conversely, $\widehat{A}_{pilot}$ shrinks as the confidence parameter $\alpha$ and variance estimate $\widehat{\Sigma}_{pilot}$ increase, reflecting greater uncertainty and a more conservative approach to covariate imbalances. Our next result shows that rerandomization with acceptance region $\widehat{A}_{pilot}$ controls the variance of the residual imbalance $R_{A}=\gamma_{0}^{\prime}Z_{h}|Z_{h}\in\widehat{A}_{pilot}$ with high probability, marginally over the realizations of the pilot data. The result is an immediate consequence of Theorem 3.5 and Theorem 5.5.
@@
Formally, the pilot estimate of $\gamma_{0}$ and Wald region could be constructed as in Robinson, (1988). A simpler practical approach suggested by the theory is to let $\widehat{\gamma}_{pilot},\widehat{\Sigma}_{pilot}$ be point and variance estimators from the regression $Y_{T}\sim 1+h+\psi$, for the “tyranny of the minority” (Lin, (2013)) outcomes $Y_{T}=(1-p)DY/p+p(1-D)Y/(1-p)$, noting that $E[Y_{T}|W]=(1-p)Y(1)+pY(0)=\bar{Y}$.